# AstroPix: A Pixelated HVCMOS Sensor for Space-Based Gamma-Ray Measurement

### v1

@@ Abstract
A next-generation medium-energy gamma-ray telescope targeting the MeV range would address open questions in astrophysics regarding how extreme conditions accelerate cosmic-ray particles, produce relativistic jet outflows, and more. One concept, AMEGO-X, relies upon the mission-enabling CMOS Monolithic Active Pixel Sensor silicon chip AstroPix. AstroPix is designed for space-based use, featuring low noise, low power consumption, and high scalability. Desired performance of the device include an energy resolution of keV (or 10% FWHM) at 122 keV and a dynamic range per-pixel of keV, enabled by the addition of a high-voltage bias to each pixel which supports a depletion depth of m. This work reports on the status of the AstroPix development process with emphasis on the current version under test, version three (v3), and highlights of version two (v2). Version 3 achieves energy resolution of at 59.5 keV and m depletion in a low-resistivity test silicon substrate.
@@ 1 Introduction
Over the past several decades, silicon strip detectors (SSDs) have been a key detector technology used in gamma-ray and cosmic-ray telescopes such as the Large Area Telescope (LAT) [1], the Alpha Magnetic Spectrometer (AMS) [2], and DArk Matter Particle Explorer (DAMPE) [3]. Breakthroughs in particle physics instrumentation have enabled the development of High Voltage-CMOS (HVCMOS) monolithic active pixel sensors (MAPS) [4], which have significant advantages over other silicon-based detectors (such as SSDs). MAPS have signal amplification and readout circuitsembedded in the sensor, enabling reductions in pixel power consumption without the need for a separate readout Application Specific Integrated Circuit (ASIC). The design reduces the overall mass of the detector system and limits the passive material in the active detector volume while also improving spatial resolution.
In particular for soft gamma-rays ( MeV), it is imperative to have two-dimensional hit location information in a single detector layer [5]. These detector capabilities make them a particularly compelling technology for future gamma-ray telescopes (see Refs. 6, 5 for more details).
AstroPix is a HVCMOS MAPS being developed for space-based mission concepts such as the All-sky Medium Energy Gamma-ray Observatory eXporer (AMEGO-X) [5]. Inspiration is drawn from the ATLASPix chip, which was designed to detect charged particles in the inner detector of the ATLAS Experiment [7]. The AstroPix project has subsequently coordinated incremental development away from the ATLASPix designs toward a final version which will be optimized for a space environment (see Table 1). The current device under test and subject of this work is the third iteration, or AstroPix_v3. The performance of AstroPix_v1 and AstroPix_v2 has been documented in Refs. 8, 9, 10, 11.
In addition to reviewing AstroPix_v3 operation and functionality, this work will overview noise and energy resolution performance. Current-voltage and capacitance-voltage curves are presented with discussion of sensor noise and depletion. AstroPix_v3 will be the first flight-tested AstroPix chip as the main component of a sounding rocket paylod, the Astropix Sounding rocket Technology dEmonstration Payload (A-STEP) (see Sec. 7). As such, operation and charicterization results relevant for this upcoming flight will be emphasized.
@@
The paper is outlined as follows: Section 2 describes the AstroPix_v3 chips; Section 3 illustrates the benchtop test setup; Section 4 describes the amplitude and impact of noise and dark count rate; Section 5 details the AstroPix_v3 characterization, including calibration, energy resolution, and depletion depth; Section 6 summarizes radiation testing of AstroPix_v2 in a heavy ion beam; and finally Section 7 summarizes the work and outlines future outlook and applications of AstroPix.
@@ 2 AstroPix_v3
HVCMOS MAPS detectors were developed by Ivan Peric more than a decade ago [4] primarily for particle physics applications. The Karlsruhe Institute of Technology (KIT) ASIC and Detector Laboratory (ADL) has continued this work, advancing the technology forward [12, 13, 14]. ADL’s experience with chips such as MuPix and ATLASpix inform the development of AstroPix. The design evolution over the course of several iterations has culminated in the first full-scale flight prototype chip: AstroPix_v3. An illustration of key properties of each AstroPix version, including AstroPix_v3, is shown in Table 1. For the first time, AstroPix_v3 uses the full cm2 reticle and features a pixel matrix with a pixel pitch of m2. A m2 high voltage deep -well (DNW) protects the embedded CMOS circuits and creates a bias junction with the -type bulk silicon, leading to a depletion region with HV application (see Fig. 1). AstroPix_v3 uses a standard high voltage CMOS process with a deep -well and unthinned bulk silicon wafers with a thickness of 720 m.
@@
ATLASPix | AstroPix_v1 | AstroPix_v2 | AstroPix_v3 | AstroPix | |
(Measured) [6, 15] | [8] | [8, 10] | [11, 16, 17] | (Goal)[5] | |
Eres (FWHM) | % | 10% | |||
at 30.1 keV | at 30.1 keV | at 30.1 keV | at 59.5 keV | at 122 keV | |
Pixel pitch [m] | |||||
Thickness [m] | 100 | 725 | 725 | 725 | 525 |
Depletion [m] | 48■ | Not measured | Not measured | 500 | |
Dynamic range | 5-32 | 14-122 | 14-80 | 14-200 | 25-700 |
[keV] | |||||
Analog power | 120 | 14.7 | 3.4 | 1.06 | 1.0 |
[mW/cm2] | |||||
Digital power | 40 | 9.9 | 3.75 | 3.06 | 0.5 |
[mW/cm2] |
@@
The signal path is shown in Fig. 2. Each pixel contains a charge-sensitive amplifier (CSA) and comparator. For testing and comparison purposes, the first three columns feature a PMOS amplifier whereas the rest implement the standard NMOS amplifiers. These PMOS columns were found to impair data quality with higher noise rates than NMOS columns, and are not considered in this work’s analysis or for future design iterations. A simplified schematic for pixel operation and timestamp generation is shown in Fig. 3.
@@
The in-pixel comparator enables self-triggering. A Time Over Threshold (ToT) measurement gives proxy for the deposited charge by recording the duration of the signal amplitude while over a user-defined threshold. In AstroPix_v3, a global comparator threshold value is set for the full array 111AstroPix_v4 and subsequent versions include Tune DACs to enable individual pixel threshold setting [18]..
@@
When a pixel comparator passes a signal that exceeds the threshold, the chip processes this information and lowers an ‘interrupt’ signal. To keep the number of readout channels low, and therefore reduce power consumption, the pixel comparator outputs per row and column are OR wired (Fig. 2). A low ‘interrupt’ signal indicates that data is ready for collection222AstroPix_v3 readout utilizes a Field Programmable Gate Array (FPGA) which receives and stores data by the software-based data acquisition (DAQ) system (see Sec. 3) The comparator threshold is set to achieve maximal detector efficiency by setting the lowest possible detector threshold while still minimizing triggering from noise fluctuations.
@@
The matrix digitisation is done via 12 bit counters driven by a MHz clock, implemented both per row and column to measure ToT (Fig. 2). Resulting ToT values fall within the s range, therefore the ns-scale resolution from this clock is much smaller than the ToT noise. An 8 bit time of arrival (ToA) timestamp is driven by a MHz clock (Fig. 2).
@@
In addition to digitized output, an “analog" output can also be accessed. This is sent from the in-pixel amplifier prior to the standard digitization readout path (including the in-pixel comparator) and can be read from a single pixel at a time from the bottom row of pixels (row0). The analog output signal provides an important cross-check on the energy resolution of the detector and front-end amplifier because it is not limited by the digital resolution of the readout.
@@
Figure 4 (left) shows a single AstroPix_v3 mounted to a carrier board. Chips are diced from the wafer as single arrays and in a -array configuration called a quad chip, shown in Fig. 4 (right) with a total area of cm2. This quad chip, utilizing four independent AstroPix arrays, will be the building block of larger AstroPix-based structures including A-STEP. of fabricated AstroPix_v3 chips (180 chips) underwent Quality Control testing following fabrication, dicing from the wafer, mounting to a carrier board, and wirebonding. of cm resistivity (Okmetic) chips passed the Quality Control testing, with the few failures due mostly to low breakdown voltage. This exemplifies the scalability that AstroPix is designed for, where the formation of large format structures with straightforward integration, operation, and readout is achievable. Power is delivered to each chip in a quad-chip via custom-designed bus bars to supply the chips in parallel. Data are sent between arrays via wirebonds that enable a daisy chain and is delivered off-chip with a quasi-Serial Peripheral Interface (SPI). Quad chips of AstroPix_v3 are also under test with a suite of software and firmware developed for the readout of large arrays and multiple layers 333https://github.com/AstroPix/astep-fw.
The MHz SPI interface clock transmits bytes in ns over two MISO lines, leading to a single chip latency of ns. Larger instrument designs can daisy chain up to AstroPix chips to one SPI bus, limited by available chip identification bits, which increases the per-bus latency up to s. The MHz timestamp clock turns over after s, so the latency does not impact timing resolution.
All iterations of AstroPix were fabricated using TSI Semiconductors’ nm process and AstroPix_v3 was delivered in October 2022 444TSI Semiconductors has subsequently been acquired by Bosch Semiconductor LLC and the foundry business has ceased. Subsequent iterations of AstroPix will be fabricated at a different foundry such as AMS in Austria or LFoundry in Italy.. The final design must deplete m and thus a global high-voltage (HV) bias is applied to a substrate contact along the chip edge on the frontside (Fig. 1). A backside bias is not utilized in order to simplify chip processing and eventual integration into large format structures. Full depletion will be achieved with a high-resistivity (cm) silicon wafer thinned to m. Neighboring pixels are isolated by a -stop isolation ring (blue in Fig. 1), separated by a distance of m. This large separation between the DNW guard ring (dark green in Fig. 1) and the isolation -stop allows for operation at high voltages with breakdown occurring at V. The bulk substrate between implants sustains the high voltage, creating a smooth depletion layer through the bulk without introducing inter-pixel dead space.
AstroPix_v3 was fabricated on , , and cm wafers provided by the TSI Foundry, Okmetic, and Topsil respectively. The highest-resistivity AstroPix_v3 chip was targeted for exploring the goal depletion depth, but the medium-resistivity chip (henseforth referred to by its manufacturer, Okmetic) will be the focus of this paper (see Sec. 5.2 for more details). The depletion value stated in Table 1 reflects this Okmetic wafer where full depletion is not expected. Further details of AstroPix characterization can be found in Refs. 11, 16, 17.
A current-voltage (IV) curve of the biasing HV used to deplete the substrate for a representative Okmetic chip is shown in Fig. 5. Breakdown occurs between V and V depending on the wafer555Breakdown occurs around V with the cm wafer.. Applications of AstroPix require low power draw and a bias leakage current less than A, allowing operation of an Okmetic AstroPix_v3 chip up to V. A larger bias enables more complete depletion (more details in Sec. 5.2), so the energy calibration reported in Sec. 5.1 uses a bias of -350 V which incorporates a safety margin with respect to the breakdown voltage. Otherwise, a stable bias of V can be assumed for the remainder of the studies in this paper unless explicitly stated.
3 Experimental Setup
@@
The equipment used to test AstroPix, as shown in Fig. 6, includes a custom built GEneric Configuration and COntrol (GECCO [19]) Data Acquisition System, a NexysVideo (Xilinx Artix-7) FPGA, and a carrier board for the integration of the chips into the GECCO system (along with an oscilloscope and power supplies). Data is read off the sensor by a software-based DAQ system.
@@
In-pixel circuitry operation can be fine-tuned through user-definable values configured through Digital-to-Analog Converters (DACs). Currents such as that fed to the source follower which sets the amplifier operating point, and voltages such as comparator voltage baseline impact signal shaping and power consumption. Initial values are set from simulation and were further optimized for performance while on the bench. Shaping the pulse width via the in-pixel band pass filter was a focus, to not limit the dynamic range given by the maximum measureable value of s by the bit MHz ToT counter. Following optimization, all testing utilizes the same set of DAC values. This can be assumed for all the results explored in this work.
@@
The chip can operate in “high gain" or “high dynamic range" mode, which impacts the in-pixel CSA gain. The studies in this work all operate in “high dynamic range" mode which causes bilinear output from the CSA. Therefore, a linear response through the full dynamic range is not expected.
@@
Custom firmware 666https://github.com/AstroPix/astropix-fw and software 777https://github.com/AstroPix/astropix-python allow for interfacing with the chip, properly setting clocks and configuring chip operational settings, and executing data collection and storage.
@@
All studies in this work exclusively consider this digital data. The OR wiring results in separate hits containing row and column information. Events are considered for data analysis only if a pair of row and column hits record a timestamp within 1 clock count ( ns) and ToT values within s. In this way, the response of individual pixels can be considered.
@@
Wafer name | Resistivity | Comparator | Bias | Max. dark |
[cm] | Thresh. [mV] | [-V] | count rate [Hz] | |
Okmetic | 200-400 | 200 | 150 | 2 |

### v2

@@ Abstract
A next-generation medium-energy gamma-ray telescope targeting the MeV range would address open questions in astrophysics regarding how extreme conditions accelerate cosmic-ray particles, produce relativistic jet outflows, and more. One concept, AMEGO-X, relies upon the mission-enabling CMOS Monolithic Active Pixel Sensor silicon chip AstroPix. AstroPix is designed for space-based use, featuring low noise, low power consumption, and high scalability. Desired performance of the device include an energy resolution of keV (or 10% FWHM) at 122 keV and a dynamic range per-pixel of keV, enabled by the addition of a high-voltage bias to each pixel which supports a depletion depth of m. This work reports on the status of the AstroPix development process with emphasis on the current version under test, version three (v3), and highlights of version two (v2). Version 3 achieves energy resolution of at keV and m depletion in a low-resistivity test silicon substrate.
keywords:
[…]
AstroPix pixelated silicon chip is being designed for use in space-based observation
AstroPix_v3 achieves energy resolution goals
First flight demonstration will feature AstroPix_v3
Continued AstroPix development will decrease power consumption and increase dynamic range
@@ 1 Introduction
Over the past several decades, silicon strip detectors (SSDs) have been a key detector technology used in gamma-ray and cosmic-ray telescopes such as the Large Area Telescope (LAT) [1], the Alpha Magnetic Spectrometer (AMS) [2], and DArk Matter Particle Explorer (DAMPE) [3]. Breakthroughs in particle physics instrumentation have enabled the development of High Voltage-CMOS (HVCMOS) Monolithic Active Pixel Sensors (MAPS) [4], which have the potential to offer significant advantages over other silicon-based detectors (such as SSDs).
MAPS have signal amplification and readout circuits embedded in the sensor, without the need for a separate Application Specific Integrated Circuit (ASIC), providing a more compact and scalable design. MAPS are traditionally low power devices, and they have by design a reduced overall mass with limited passive material in the active area of the detector. The two-dimensional hit location information provided by these pixelated detectors is an imperative feature for soft gamma-ray detection ( MeV). These detector capabilities make them a particularly compelling technology for future gamma-ray telescopes (see Refs. 5, 6 for more details).
AstroPix is a HVCMOS MAPS being developed for space-based mission concepts such as the All-sky Medium Energy Gamma-ray Observatory eXplorer (AMEGO-X) [6]. Inspiration is drawn from the ATLASPix chip, which was designed to detect charged particles in the inner detector of the ATLAS Experiment [7]. The AstroPix project has subsequently coordinated incremental development away from the ATLASPix designs toward a final version which will be optimized for a space environment (see Table 1). The device under test and subject of this work is the third iteration, or AstroPix_v3. The performance of AstroPix_v1 and AstroPix_v2 has been documented in Refs. 8, 9, 10, 11.
In addition to reviewing AstroPix_v3 operation and functionality, this work will overview noise and energy resolution performance. Current-voltage and capacitance-voltage curves are presented with discussion of sensor noise and depletion. AstroPix_v3 will be the first flight-tested AstroPix chip as the main component of a sounding rocket payload, the Astropix Sounding rocket Technology dEmonstration Payload (A-STEP) (see Sec. 7). As such, operation and characterization results relevant for this upcoming flight will be emphasized.
@@
The paper is outlined as follows: Section 2 describes the AstroPix_v3 chips; Section 3 illustrates the benchtop test setup; Section 4 describes measurements and impact of noise and dark count rate; Section 5 details the AstroPix_v3 characterization, including calibration, energy resolution, and depletion depth; Section 6 summarizes radiation testing of AstroPix_v2 in a heavy ion beam; and finally Section 7 summarizes the work and outlines future outlook and applications of AstroPix.
@@ 2 AstroPix_v3
HVCMOS MAPS detectors were developed by I. Peric more than a decade ago [4] as a novel technology primarily for particle physics applications. His group at the Karlsruhe Institute of Technology (KIT) ASIC and Detector Laboratory (ADL), with many more collaborators, has continued this work, advancing the technology forward [12, 13, 14]. ADL’s experience with chips such as MuPix and ATLASpix inform the development of AstroPix. The design evolution over the course of several iterations has culminated in the first full-scale flight prototype chip: AstroPix_v3. An illustration of key properties of each AstroPix version is shown in Table 1. For the first time, AstroPix_v3 uses the full cm2 reticle and features a pixel matrix with a pixel pitch of m2. A m2 high voltage deep -well (DNW) protects the embedded CMOS circuits and creates a bias junction with the -type bulk silicon, leading to a depletion region with high voltage (HV) application (see Fig. 1). HV is applied as a negative voltage to the p+ ring near the chip edge. In-pixel collection diodes are biased to V. Fabrication uses a standard high voltage CMOS process with a deep -well and unthinned bulk silicon wafers with a thickness of 720 m.
| ATLASPix | v1 | v2 | v3 | Goal | |
| [5, 15] | [8] | [8, 10] | [11, 16, 17] | [6] | |
| Eres [keV] | % | 10% | |||
| (FWHM) | at 30.1 | at 30.1 | at 30.1 | at 59.5 | at 122 |
| Pitch [m] | |||||
| Thick [m] | 100 | 725 | 725 | 725 | 525 |
| Dep. [m] | 48■ | N.M. | N.M. | 500 | |
| Range [keV] | 5-32 | 14-122 | 14-80 | 22-200 | 25-700 |
| Analog | 120 | 14.7 | 3.4 | 1.06 | 1.0 |
| Digital | 40 | 9.9 | 3.75 | 3.06 | 0.5 |
@@
All iterations of AstroPix were fabricated using TSI Semiconductors’ nm process 111TSI Semiconductors has subsequently been acquired by Bosch Semiconductor LLC and the foundry business has ceased. Subsequent iterations of AstroPix will be fabricated at a different foundry such as AMS in Austria or LFoundry in Italy. and AstroPix_v3 was delivered in October 2022. The final design must deplete m and thus a global HV bias is applied to the chip top side (Fig. 1). A backside bias is not utilized in order to simplify chip processing and eventual integration into large format structures. Full depletion will be achieved with a high-resistivity ( k) silicon wafer thinned to m. Neighboring pixels are isolated by a -stop isolation ring (blue in Fig. 1), separated by a distance of m. The metal DNW guard ring (dark green in Fig. 1) is connected to each pixel DNW and overhags the DNW and silicon bulk by m on each side. It acts as a field plate to reduce the electrical field strength at the implant edges, thus improving breakdown voltage. The large separation between the DNW guard ring and the isolation -stop allows for operation at high voltages with breakdown occurring at V. The bulk substrate between implants sustains the high voltage, creating a smooth depletion layer through the bulk without introducing inter-pixel dead space. The m distance between edge pixels and the chip edge, as well as the final p-well block, is to ensure that the depletion region does not touch the chip edge.
Chips were fabricated on three different wafers with low, medium, and high resistivity (see Table 2). The medium-resistivity Wafer 2 is the emphasis of this work. Further details of AstroPix characterization can be found in Refs. 11, 16, 17.
| Wafer | 1 | 2 | 3 |
| Resistivity [] | 20 | 200-400 | 20900-28900 |
| Manufacturer | TSI | Okmetic | Topsil |
| Notes | Fabricator test wafer | Focus of this work, | Tested in this work |
| with poorly quantified | used to populate Table 1, | to explore the goal | |
| resistivity | full depletion not expected | depletion depth |
A current-voltage (IV) curve of the biasing HV used to deplete the substrate for a representative wafer 2 chip is shown in Fig. 2. The DNW guard rings are electrically coupled to the DNW for this measurement. Breakdown occurs between V and V in wafer 2 and around V in wafer 1. Applications of AstroPix require low power draw and a bias leakage current less than A, allowing operation of a wafer 2 chip up to V. A larger bias enables more complete depletion (more details in Sec. 5.2), so the energy calibration reported in Sec. 5.1 uses a bias of -350 V which incorporates a safety margin with respect to the breakdown voltage. Otherwise, a stable bias of V can be assumed for the remainder of the studies in this paper unless explicitly stated.
@@
A simplified schematic for pixel operation and timestamp generation is shown in Fig. 3. Each pixel contains a charge-sensitive amplifier (CSA), bandpass filter, and comparator. For testing and comparison purposes, the first three columns feature a PMOS amplifier whereas the rest implement the standard NMOS amplifiers. These PMOS columns were found to impair data quality with higher noise rates than NMOS columns, and are not considered in this work’s analysis or for future design iterations.
@@
Each pixel contains a band pass filter to mitigate thermal and flicker noise (see Sec. 4). The in-pixel comparator (‘comp’ in Fig. 3) enables self-triggering, so no external trigger signal is required for readout and the device only returns data when an event occurs. This reduces the data footprint as pixels with no measured charge do not return zero or empty counts. A Time Over Threshold (ToT) measurement provides a measurement of the deposited charge by recording the duration of the signal amplitude while it remains above a user-defined threshold. The timing of this signal is set by an external global timestamp (TS).In AstroPix_v3, a global comparator threshold value is set for the full array. AstroPix_v4 and subsequent versions include Tune DACs to enable individual pixel threshold setting [18].
@@
The full signal path is shown in Fig. 4. When a pixel comparator passes a signal that exceeds the threshold, the chip processes this information and lowers an ‘interrupt’ signal. To keep the number of readout channels low, and therefore reduce power consumption, the pixel comparator outputs per row and column are OR wired (Fig. 4). A low ‘interrupt’ signal indicates that data is ready for collection to a Field Programmable Gate Array (FPGA) which receives and stores data then transmitted to a software-based data acquisition (DAQ) system (see Sec. 3). The comparator threshold is set to achieve maximal detector efficiency by setting the lowest possible detector threshold while still minimizing triggering from noise fluctuations.
@@
The matrix digitization is done via 12 bit counters driven by a MHz ToT clock, implemented both per row and per column (Fig. 4). Resulting ToT values fall within the s range, therefore the ns-scale resolution from this clock is much smaller than the ToT noise. An 8 bit time of arrival (ToA) timestamp is driven by a MHz clock.
@@
In addition to digitized output, an “analog” output can also be accessed. This is sent from the in-pixel amplifier prior to the standard digitization readout path (including the in-pixel comparator) and can be read from a single pixel at a time from the bottom row of pixels (row0). The analog output signal provides an important cross-check on the energy resolution of the detector and front-end amplifier because it is not limited by the digital resolution of the readout.
@@
Figure 5 (left) shows a single AstroPix_v3 chip mounted to a carrier board. Chips are diced from the wafer as single arrays and in a -array configuration called a quad chip, shown in Fig. 5 (right) with a total area of cm2. This quad chip, utilizing four independent AstroPix arrays, will be the building block of larger AstroPix-based structures including A-STEP [19]. of fabricated chips (180 chips) underwent Quality Control testing following fabrication, dicing from the wafer, mounting to a carrier board, and wirebonding. Individual chips were tested with an IV measurement, visual inspection for physical damage, and metrology for dimension and mass. Passing chips displayed no physical damage inside guard rings and a sufficiently high breakdown voltage around V. of wafer 2 chips passed the Quality Control testing, with the few failures due mostly to low breakdown voltage. No ASIC readout testing was conducted at this stage, however this is planned in the future as active work is underway to commission the necessary infrastructure and develop a test procedure. This chip yield exemplifies the scalability that AstroPix is designed for, where the formation of large format structures with straightforward integration, operation, and readout is achievable. Power is delivered to each chip in a quad-chip via custom-designed bus bars to supply the chips in parallel. Data are sent between arrays via wirebonds that enable a daisy chain and is delivered off-chip with a quasi-Serial Peripheral Interface (SPI). Quad chips of AstroPix_v3 are also under test with a suite of software and firmware developed for the readout of large arrays and multiple layers 222https://github.com/AstroPix/astep-fw. Larger instrument designs can daisy chain up to chips to one SPI bus, limited by available chip identification bits. The chip-to-chip transfer latency is expected to be ns total with ns/byte and assuming idle bytes per hop. The MHz timestamp clock turns over after s, so the latency does not impact timing resolution.
3 Experimental Setup and Sensor Operation
@@
The equipment used to test AstroPix, as shown in Fig. 6, includes a custom built GEneric Configuration and COntrol (GECCO [20]) Data Acquisition System, a NexysVideo (Xilinx Artix-7) FPGA, and a carrier board for the integration of the chips into the GECCO system, along with an oscilloscope and power supplies. Data is read off the sensor by a software-based DAQ system.
@@
In-pixel circuitry operation can be fine-tuned through user-definable values configured through in-chip Digital-to-Analog Converters (DACs). Currents, such as that fed to the source follower which sets the amplifier operating point, and voltages, such as comparator voltage baseline, impact signal shaping and power consumption. Initial values are set from simulation and were further optimized for performance while on the bench. Shaping the pulse width via the in-pixel band pass filter was a focus of optimization studies, to not limit the operational range given by the maximum measurable value of s by the bit MHz ToT counter. Following optimization, all testing utilizes the same set of DAC values. This can be assumed for all the results explored in this work.
@@
The chip can operate in “high gain” or “high dynamic range” mode, which impacts the in-pixel CSA gain. The studies in this work all operate in “high dynamic range” mode which causes bilinear output from the CSA. Therefore, a linear response through the full operational range is not expected.
@@
Custom firmware 333https://github.com/AstroPix/astropix-fw and software 444https://github.com/AstroPix/astropix-python allow for interfacing with the chip, properly setting clocks and configuring chip operational settings, and executing data collection and storage.
@@
All studies in this work exclusively consider this digital data. The row- and column-wise OR wiring results in the data from a pixel being separated in two hits containing the row and column information. Events are considered for data analysis only if a pair of row and column hits record a timestamp within 1 clock count ( ns) and ToT values within s. In this way, the response of individual pixels can be considered.
@@
| Wafer | Wafer | Resistivity | Comparator | Bias | Max. dark |
| identifier | brand | [] | Thresh. [mV] | [-V] | count rate [Hz] |
| 2 | Okmetic | 200-400 | 200 | 150 | 2 |